OpenStudy (anonymous):

Does the function y=(x+2)/x have an inverse function? I would say no because it has a vertical asymptote at x=0 and it's not continuous, plus I can't fix it up to just one variable of y

OpenStudy (anonymous):

I turned y=x+2/x into y=(x^2+2)/x using the standard 'inversion' procedure. I graphed the 2 equations to see what I'd end up with: I got something like this.  I don't think a function and its inverse necessarily have to be continuous; all an inverse is really is a close reflection of the original equation around the origin, right? Excuse me if I'm wrong. But I'd think that y=x+2/x may have an inverse.

OpenStudy (anonymous):

Ok, but I have to write what it is, and my book doesn't just have x=(y+2)/y as an answer ever it ususally is like isolate the y, but I can't here

OpenStudy (primeralph):

Why can't you?

OpenStudy (primeralph):